We are given AB=12,CD=6,EF=18 and GH=24
Pair up the equal measures
GH+EF
EF-CD
GH-EF
AB-CD
GH-AB
EF+2AB
Please work out answers

Answer by Mathtut(3670)

I will do a few
:
GH+EF=24+18=42
:
EF-CD=18-6=12
: